OpenStudy (ujjwal):

The rms speed of molecules in a mono-atomic gas is C. The speed of sound in the gas is: a)C b)\(\frac{2}{3}\)C c)\(\frac{\sqrt5}{3}\)C d)\(\sqrt5\)C

OpenStudy (anonymous):

Ans: C

OpenStudy (anonymous):

\[V_{rms}=\sqrt{\frac{3RT}{M}}\] and \[V_{sound}=\sqrt{\frac{\gamma RT}{M}}\] where \[\gamma=1+\frac{2}{f}\] where f is the degree of freedom. Here f=3 for mono-atomic gas. Now on using above formulas, you can get the answer. If you stuck somewhere then plz inform me.

OpenStudy (ujjwal):

Dhanyabad @ramkrishna sir!